2.00 points Engberg Company installs lawn sod in home yards. The companys most recent monthly contribution formatincome Sales Variable expenses Contribution margin Fixed expenses Net operating income Amount of Sales 49,200 3,800 19,000 54,800 $ 123,000 100% 40% 60% Required: 1. Compute the companys degree of operating leverage. (Round your answer to 2 decimal places.)
8 Pek Pines Mtn Coty 38R Estes Park Required: . Compute the companys degree of operating leverage. (Round your answer to 2 decimal places.) 2 Using the dogree o, operating Round your percentage answer to 2 decimal places (o.123 leverage, estimate the impact on net operating income of a 27 % increase in sales (Round your intermediate calculations to 2 decimal place 1 3 Construct a new contribution format income statement for the company assuming a 27% increase in sale
3. Construct a new contribution format income statement for the company assuming a 2 7% increase in sales. Engberg C Contribution Income Statement Total
Answer #1

Part 1) degree of operating leverage = contribution margin/ net operating income

= $73800/$54800 = 1.35

Part 2)

Estimated percent change in net operating income =

1.3467 × 27% = 36.36%

Part 3)

Sales (123000 × 1.27) $156210
Variable expenses ( 49200 × 1.27) ($62484)
Contribution margin $93726
Fixed expense ($19000)
Net operating income $74726
Original net operating income $54800


Percent change in net operating income = ($74726-$54800)/54800 = 36.36%

I have given complete and detailed solution to your problem, in case of any query you can comment and

PLEASE DO GIVE POSITIVE RATING